OpenStudy (anonymous):

The time a cake must bake is between 25 and 30 minutes, inclusive. A) 25 < x < 30 B) 25 *less than or equal to* x *less than or equal to* 30 C) 25 *less than or equal to* x *greater than or equal to* 30 D) 25 < x > 30

OpenStudy (danjs):

let x = the time the cake bakes

OpenStudy (danjs):

the time is inbetween 25 and 30 and includes 25 and 30 exactly

OpenStudy (danjs):

< and > means the end points are not included \[\ge~~or~~\le ~~~endpoints ~are~included\]

OpenStudy (danjs):

so you have x is greater than or equal(inclusive) 25 AND x is less than or equal (inclusive) to 30 \[25 \le x \le 30\]

OpenStudy (anonymous):

So A ?

OpenStudy (littlebird):

Do you understand the meanings of the signs?

OpenStudy (anonymous):

yeah I do. I just want to make sure i got the right answer so i don't get it wrong. Thanks :)

OpenStudy (littlebird):

It is B. The problem says that the times are inclusive. You show inclusion by writing "less than or equal to" or "greater than or equal to" A does not have any of this.

OpenStudy (anonymous):

You mean type in the signs of less or greater than or equal to right?

OpenStudy (littlebird):

Yes I'm on my iPad, using safari, so I can't type the signs.

OpenStudy (anonymous):

Neither can i lol thats why i use my special *. . . *'s